# UDPNet: Unleashing Depth-based Priors for Robust Image Dehazing

### v1

@@ 6 Conclusion
In recent years, the field of single-image dehazing has witnessed significant advancements, primarily driven by effective module design. However, the utilization of multimodal information remains limited, and depth information that is closely tied to the physical process of haze formation is often overlooked. In this paper, we present UDPNet, a general framework guided by depth priors for image dehazing to enhance multi-scale, multi-shape, and dual-modal representation learning. Our approach integrates a large-scale pre-trained depth estimation model to extract scene depth priors, which are utilized within diverse U-shaped network architectures. By leveraging efficient fusion strategies, we adaptively tailor the model to handle different haze distributions. These combined innovations yield a robust solution for image dehazing, as demonstrated by our extensive evaluations, while also showcasing the potential of our framework for universal image restoration. Our future work will explore the framework in other image restoration tasks.

### v2

@@ VI Conclusion
In recent years, the field of single-image dehazing has witnessed significant advancements, primarily driven by effective module design. However, the utilization of multimodal information remains limited, and depth information that is closely tied to the physical process of haze formation is often overlooked. In this paper, we present UDPNet, a general framework guided by depth priors for image dehazing to enhance multi-scale and dual-modal representation learning. Our approach integrates a large-scale pre-trained depth estimation model to extract scene depth priors, which are utilized within diverse U-shaped network architectures. By leveraging efficient fusion strategies, we adaptively tailor the model to handle different haze distributions. These combined innovations yield a robust solution for image dehazing, as demonstrated by our extensive evaluations, while also showcasing the potential of our framework for universal image restoration.